OpenStudy (anonymous):

what is the absolute maximum for 4 cos x+2sin 2x

Parth (parthkohli):

Ooo, ooo. So many ways to do this. The easiest would be to compute the derivative and equate it to zero.

OpenStudy (anonymous):

i already did that but the answer i got wasn't the right one

Parth (parthkohli):

\[-4\sin (x) + 4\cos(2x) = 0\]

Parth (parthkohli):

That is to say, \(\cos(2x) = \sin(x)\) or \(1 - 2\sin^2(x) = \sin(x)\). That is a quadratic equation and you solve it like one.